OpenStudy (yourboydusty):

The circle below has a radius of 10 cm. What is the area of the shaded region? If necessary, round your answer to two decimal places. Do not include units in your answer.

OpenStudy (studygurl14):

First find teh area of the whole circle, then multiply it by 90/360

OpenStudy (yourboydusty):

i have like 52 questions to do and not alot of time lol

OpenStudy (yourboydusty):

um okay 1 sec

Directrix (directrix):

The area of a circle is given by A = pi * r^2 90/360 = 1/4 1/4 of the circle is shaded. Crank out this: 1/4 * 3.14 * 10 * 10 and see what you get. 3.14 is an approximation for pi.

OpenStudy (yourboydusty):

okay so i got 78.5?

Directrix (directrix):

The exact answer is 25 * pi If the approximation for pi as 3.14 is not used, then the answer to two digits becomes 78.54 Look at the instructions for these problems and see if you are told what to use as the approximation for pi.

OpenStudy (yourboydusty):

uuummm so what do i need to do srry i dont understand lel

OpenStudy (yourboydusty):

@Directrix

Directrix (directrix):

Do the instructions for this problem tell you what number to substitute for pi in the formula for the area of a circle?

Directrix (directrix):

1/4 * pi * 10 * 10 = 78.5398..... which to two digits is 78.54. Go with that if you have no instructions for what to use an an approximation for pi.